A planet has twice the mass of Earth and twice the radius of Earth. The free-fall acceleration on this planet is equal to what?
Answer #1

g is proportional to mass and inversely proportional to radius of the earth

g would be              gearth/gplanet = m/r2 * (2r)2/2m
So gplanet = 1/2 * gearth = 0.5 * 9.8 = 4.9 m/s2(ans)